# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ B.2.1 Initial preprocessing steps
In the following, we describe the parameterless and pre-specified preprocessing steps that are applied to the full COMPANION data set in its rawest version available. Note that the raw data set is on patient contact level, which was the unit for data collection (Hodiamont et al.,, 2022). The initial preprocessing steps are:
@@
    the removal of contacts with palliative care phase “bereavement”, AKPS $=0$ (“dead”), or costs $=0$,
@@ B.2.2 Preprocessing steps in the learning pipeline
The distribution of the outcome variable in the COMPANION data set is right skewed, i.e. some palliative care phases have exceptionally high costs (see Table S1). Since it is not possible to definitively attribute these values to data entry errors, they are not permanently removed from the data set. However, since the prediction values calculated by the corresponding decision tree algorithm in each terminal node can be sensitive to outliers, removing cost outliers during the training process could improve model performance. Importantly, this preprocessing step is only applied during training and not during prediction, i.e. when the final prediction model is used to make predictions on a data set, no cost outliers are removed. Removing them during prediction could artificially improve the model’s performance, as cost outliers are typically difficult to predict correctly (see also Kapoor and Narayanan,, 2023).
The definition of outliers is generally not straightforward, as many possible options exist (Kuhn and Johnson,, 2013; Steyerberg,, 2019). We denote the corresponding HP as $\lambda_{outlier}$. In our illustration, we define all cost values higher than the $\lambda_{outlier}$th cost percentile as outliers, with $\lambda_{outlier}\in\{100,99,95,90\}$. If $\lambda_{outlier}=100$ (the default value), no outliers are removed. Note that this preprocessing step includes the parameter $\theta_{outlier}$, which corresponds to the percentile calculated according to $\lambda_{outlier}$.
@@
As outlined in Section 5.1, the set of features to generate the prediction model includes the Integrated Palliative care Outcome Scale (IPOS; Murtagh et al.,, 2019), which is a score based on 17 individual features covering physical symptoms, psycho-social burden, family needs, and practical problems. Each of the 17 features is ordinal and can take values from 0 to 4, where 0 and 4 correspond to the least and highest symptom or concern severity, respectively. For example, for the features IPOS-“Pain” and IPOS-“Shortness of Breath”, a value of 0 corresponds to “not at all” and a value of 4 corresponds to “overwhelmingly” (see Figure S1 for an overview of all 17 features).
@@
While observations affected by the first type of missing values (“missing”) do not occur often and are removed as part of the initial preprocessing steps described in Section B.2.1, handling the “cannot assess” values is more challenging. If all observations with at least one “cannot assess” response were removed, almost half of the COMPANION data set would be discarded (see Table S2; this would also apply approximately to any subset $\mathcal{D}_{\text{train}}$ or $\mathcal{D_{\text{new}}}$ of the COMPANION data set). To avoid the loss of valuable information, an alternative approach is to treat “cannot assess” values as 0 (i.e. least symptom or concern severity), based on the assumption that an unobserved burden does not initiate a care mandate and therefore does not result in costs. However, it is not clear whether this assumption is valid for observations where many or even all IPOS features are recorded as “cannot assess” (e.g., if 15 out of 17 IPOS features are recorded as “cannot assess”, these features might not have been assessed at all). It could thus be a reasonable approach to set “cannot assess” values to 0 but exclude observations with many “cannot assess” values, as they potentially result in incorrect IPOS scores. Specifying the exact threshold for the maximum number of “cannot assess” values is, however, not straightforward. It can be denoted as HP $\lambda_{ca}$, and ranges from 0 to 17 (observations with more than $\lambda_{ca}$ “cannot assess” values are removed; if $\lambda_{ca}$ = 17, no observations are removed). In our illustration, we consider the values $\{16,14,12,10\}$ for $\lambda_{ca}$, with $\lambda_{ca}=16$ being the default.
@@
Instead of aggregating all 17 IPOS features into one score, it is also possible to generate multiple IPOS scores based on the subscales in which the features can be divided (Murtagh et al.,, 2019). These subscales are: (i) physical symptoms (10 features), (ii) emotional symptoms (4 features), and (iii) communication/practical issues (3 features) (see Figure S1). In our illustration, we consider the generation of two subscale scores: one score that sums the features corresponding to the physical symptoms (IPOS-physical; $[0,40]$) and one score that sums the remaining features (IPOS-others; $[0,28]$). Note that in this case, the number of features provided to the learning algorithm increases from $p=6$ to $p=7$.
@@
In the COMPANION data set, age is measured on an integer scale and ranges from 23 to 102 years (see Table S1). In its default configuration, this feature is provided to the learning algorithm on its original integer scale, without any preprocessing. Alternatively, it could be transformed into a categorical feature with six categories, using the years 50, 60, 70, 80, and 90 as cutpoints. This option could improve the model’s prediction error, as, for example, the CART algorithm suffers from a selection bias towards features with many possible splits (Hothorn et al.,, 2006). We refer to the HP that specifies the used option as $\lambda_{age}$, with no modification of age as default. This preprocessing step has the same characteristics as the aggregation of individual IPOS features into a score (i.e. no parameters, applied during training and prediction).
@@
The Australian Karnofsky Performance Score (AKPS; Abernethy et al.,, 2005), which measures patients’ functional status on an ordinal scale, takes values of $\{10,20,...,90\}$ in the COMPANION data set, with AKPS $=10$ corresponding to “comatose or barely rousable” and AKPS $=90$ to “able to carry on normal activity; minor sign of symptoms of disease” (see Table S1). In its default configuration, AKPS is considered ordinal, with the three highest categories, 70, 80, and 90, merged due to their low frequency. However, it might also be reasonable to transform AKPS into an unordered categorical variable, as costs may not monotonically decrease or increase with AKPS, but could be highest when the patient has, for example, an AKPS of 50, which corresponds to “considerable assistance and frequent medical care required”. In this case, we collapse the AKPS categories even further to avoid overfitting, resulting in AKPS $\in\{10\textnormal{-}20,30\textnormal{-}50,60\textnormal{-}90\}$. We refer to the corresponding HP as $\lambda_{akps}$, with the ordered AKPS variable as default. This preprocessing step has the same characteristics as the two previous preprocessing steps (i.e. no parameters, applied during training and prediction).
@@ B.3 Performance measures
In the illustration, two performance measures are considered: RMSE and $R^{2}$. The RMSE is obtained by taking the square root of the MSE (see Section 3.1) and is expressed in the same units as the outcome variable (i.e. costs in €). It ranges from 0 to $\infty$, where $\text{RMSE}=0$ indicates perfect prediction. The $R^{2}$ performance measure is calculated by dividing the squared error of the prediction model by the squared error of a naive model that predicts the mean and then subtracting this ratio from 1. It is a relative measure that can be interpreted as the proportion of variance in the outcome variable explained by the prediction model. The range of $R^{2}$ is $(-\infty,1]$, with $R^{2}=1$ indicating perfect prediction and a $R^{2}$ value of 0 or less indicating that a model performs no better or worse than the naive model, respectively. In this context, a lower prediction error corresponds to a higher $R^{2}$ value. See, e.g., Kuhn and Johnson, (2013) for more details on both performance measures.
@@ B.5 Clustering structure
In Figure 5 (Section 5.3), which presents the prediction error differences for 96 analysis settings, it can be seen that the CV error unexpectedly exhibits an optimistic bias in settings without HP tuning. The same observation applies to the nested CV error in analysis settings with HP tuning. These results can be attributed to the clustering structure of the COMPANION data set, and we will explain this in more detail below. Specifically, we will describe the clustering structure (Section B.5.1), explain how it impacts the estimated prediction errors (Section B.5.2), discuss why the experimental setup was not adapted to account for this clustering (Section B.5.3), and present an additional extension of the experimental setup with respect to clustering (Section B.5.4).
@@ B.5.2 Impact on prediction error estimates
For the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$, which is the focus of our illustration, this has two key implications: If $\widehat{\mathrm{PE}}_{\text{train}}$ results from an analysis setting where the apparent error was used to evaluate the final prediction model, the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ may underestimate the optimistic bias that would arise if $\mathcal{D_{\text{new}}}$ contained exclusively observations from new episodes and teams not present in $\mathcal{D}_{\text{train}}$. If $\widehat{\mathrm{PE}}_{\text{train}}$ corresponds to the simple or nested CV error, the clustering-induced optimistic bias would – under the assumption that $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ are subject to the same level of bias – effectively cancel out when considering the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$. However, as shown in Figure 5, this is not the case. Further analysis (not shown) reveals that the observed differences arise from the slightly higher proportion of patient episodes present in both $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ during resampling, compared to the proportion of episodes present in both $\mathcal{D}_{\text{train}}$ and $\mathcal{D_{\text{new}}}$ during the initial split. As a result, $\widehat{\mathrm{PE}}_{\text{train}}$ is affected by a larger optimistic bias than $\widehat{\mathrm{PE}}_{\text{new}}$, which manifests in Figure 5, where their difference is examined.
@@ B.5.4 Learning algorithms for clustered data
In addition to performing splits at the cluster level, we also extended the main experimental setup by including additional learning algorithms specifically designed for clustered data. These are the Random Effects/Expectation-Maximization Tree algorithm (REEMT; R package REEMtree; Sela and Simonoff,, 2011), and the Linear Mixed-Effects Model Tree algorithm (LMMT; R package glmertree; Fokkema et al.,, 2017). In the implementation used for our illustration, both algorithms take into account the clustering structure by iterating between two steps: (i) fitting a decision tree using the CART algorithm for REEMT or the CIT algorithm for LMMT and (ii) estimating random intercepts via a linear mixed model, which are subtracted from the outcome variable in the subsequent tree-fitting iteration. To ensure model stability, random effects are only included for each palliative care team, rather than for each individual episode, as more than 300 episodes consist of only a single palliative care phase (Figure S7a). Including REEMT and LMMT in the analysis, however, does not yield new insights. Their results closely resemble those of CART and CIT, as demonstrated in Figure S9, which compares the prediction error differences of the algorithms.

### v2

@@ B.2.1 Initial preprocessing steps
In the following, we describe the parameterless and pre-specified preprocessing steps that are applied to the full COMPANION data set in its rawest version available. Note that the raw data set is on patient contact level, which was the unit for data collection [99]. The initial preprocessing steps are:
@@
    the removal of contacts with palliative care phase “bereavement”, AKPS =0=0 (“dead”), or costs =0=0,
@@ B.2.2 Preprocessing steps in the learning pipeline
The distribution of the outcome variable in the COMPANION data set is right skewed, i.e.  some palliative care phases have exceptionally high costs (see Table S1). Since it is not possible to definitively attribute these values to data entry errors, they are not permanently removed from the data set. However, since the prediction values calculated by the corresponding decision tree algorithm in each terminal node can be sensitive to outliers, removing cost outliers during the training process could improve model performance. Importantly, this preprocessing step is only applied during training and not during prediction, i.e. when the final prediction model is used to make predictions on a data set, no cost outliers are removed. Removing them during prediction could artificially improve the model’s performance, as cost outliers are typically difficult to predict correctly [108, see also].
The definition of outliers is generally not straightforward, as many possible options exist [111, 132]. We denote the corresponding HP as $\lambda_{outlier}$. In our illustration, we define all cost values higher than the $\lambda_{outlier}$th cost percentile as outliers, with $\lambda_{outlier}\in\{100,99,95,90\}$. If $\lambda_{outlier}=100$ (the default value), no outliers are removed. Note that this preprocessing step includes the parameter $\theta_{outlier}$, which corresponds to the percentile calculated according to $\lambda_{outlier}$.
@@
As outlined in Section 5.1, the set of features to generate the prediction model includes the Integrated Palliative care Outcome Scale (IPOS; [117]), which is a score based on 17 individual features covering physical symptoms, psycho-social burden, family needs, and practical problems. Each of the 17 features is ordinal and can take values from 0 to 4, where 0 and 4 correspond to the least and highest symptom or concern severity, respectively. For example, for the features IPOS-“Pain” and IPOS-“Shortness of Breath”, a value of 0 corresponds to “not at all” and a value of 4 corresponds to “overwhelmingly” (see Figure S1 for an overview of all 17 features).
@@
While observations affected by the first type of missing values (“missing”) do not occur often and are removed as part of the initial preprocessing steps described in Supplementary Section B.2.1, handling the “cannot assess” values is more challenging. If all observations with at least one “cannot assess” response were removed, almost half of the COMPANION data set would be discarded (see Table S2; this would also apply approximately to any subset $\mathcal{D}_{\text{train}}$ or $\mathcal{D_{\text{new}}}$ of the COMPANION data set). To avoid the loss of valuable information, an alternative approach is to treat “cannot assess” values as 0 (i.e.  least symptom or concern severity), based on the assumption that an unobserved burden does not initiate a care mandate and therefore does not result in costs. However, it is not clear whether this assumption is valid for observations where many or even all IPOS features are recorded as “cannot assess” (e.g., if 15 out of 17 IPOS features are recorded as “cannot assess”, these features might not have been assessed at all). It could thus be a reasonable approach to set “cannot assess” values to 0 but exclude observations with many “cannot assess” values, as they potentially result in incorrect IPOS scores. Specifying the exact threshold for the maximum number of “cannot assess” values is, however, not straightforward. It can be denoted as HP $\lambda_{ca}$, and ranges from 0 to 17 (observations with more than $\lambda_{ca}$ “cannot assess” values are removed; if $\lambda_{ca}$ = 17, no observations are removed). In our illustration, we consider the values $\{16,14,12,10\}$ for $\lambda_{ca}$, with $\lambda_{ca}=16$ being the default.
@@
Instead of aggregating all 17 IPOS features into one score, it is also possible to generate multiple IPOS scores based on the subscales in which the features can be divided [117]. These subscales are: (i) physical symptoms (10 features), (ii) emotional symptoms (4 features), and (iii) communication/practical issues (3 features) (see Figure S1). In our illustration, we consider the generation of two subscale scores: one score that sums the features corresponding to the physical symptoms (IPOS-physical; $[0,40]$) and one score that sums the remaining features (IPOS-others; $[0,28]$). Note that in this case, the number of features provided to the learning algorithm increases from $p=6$ to $p=7$.
@@
In the COMPANION data set, age is measured on an integer scale and ranges from 23 to 102 years (see Table S1). In its default configuration, this feature is provided to the learning algorithm on its original integer scale, without any preprocessing. Alternatively, it could be transformed into a categorical feature with six categories, using the years 50, 60, 70, 80, and 90 as cutpoints. This option could improve the model’s prediction error, as, for example, the CART algorithm suffers from a selection bias towards features with many possible splits [105]. We refer to the HP that specifies the used option as $\lambda_{age}$, with no modification of age as default. This preprocessing step has the same characteristics as the aggregation of individual IPOS features into a score (i.e. no parameters, applied during training and prediction).
@@
The Australia-modified Karnofsky Performance Status (AKPS; [72]), which measures patients’ functional status on an ordinal scale, takes values of $\{10,20,...,90\}$ in the COMPANION data set, with AKPS =10=10 corresponding to “comatose or barely rousable” and AKPS =90=90 to “able to carry on normal activity; minor sign of symptoms of disease” (see Table S1). In its default configuration, AKPS is considered ordinal, with the three highest categories, 70, 80, and 90, merged due to their low frequency. However, it might also be reasonable to transform AKPS into an unordered categorical variable, as costs may not monotonically decrease or increase with AKPS, but could be highest when the patient has, for example, an AKPS of 50, which corresponds to “considerable assistance and frequent medical care required”. In this case, we collapse the AKPS categories even further to avoid overfitting, resulting in AKPS $\in\{10\textnormal{-}20,30\textnormal{-}50,60\textnormal{-}90\}$. We refer to the corresponding HP as $\lambda_{akps}$, with the ordered AKPS variable as default. This preprocessing step has the same characteristics as the two previous preprocessing steps (i.e. no parameters, applied during training and prediction).
@@ B.3 Performance measures
In the illustration, two performance measures are considered: RMSE and $R^{2}$. The RMSE is obtained by taking the square root of the MSE (see Section 3.1) and is expressed in the same units as the outcome variable (i.e. costs in €). It ranges from 0 to $\infty$, where $\text{RMSE}=0$ indicates perfect prediction. The $R^{2}$ performance measure is calculated by dividing the squared error of the prediction model by the squared error of a naive model that predicts the mean and then subtracting this ratio from 1. It is a relative measure that can be interpreted as the proportion of variance in the outcome variable explained by the prediction model. The range of $R^{2}$ is $(-\infty,1]$, with $R^{2}=1$ indicating perfect prediction and a $R^{2}$ value of 0 or less indicating that a model performs no better or worse than the naive model, respectively. In this context, a lower prediction error corresponds to a higher $R^{2}$ value. See, e.g., [111] for more details on both performance measures.
@@ B.5 Clustering structure
In Figure 5 (Section 5.3), which presents the prediction error differences for 96 analysis settings, it can be seen that the CV error unexpectedly exhibits an optimistic bias in settings without HP tuning. The same observation applies to the nested CV error in analysis settings with HP tuning. These results can be attributed to the clustering structure of the COMPANION data set, and we will explain this in more detail below. Specifically, we will describe the clustering structure (Supplementary Section B.5.1), explain how it impacts the estimated prediction errors (Supplementary Section B.5.2), discuss why the experimental setup was not adapted to account for this clustering (Supplementary Section B.5.3), and present an additional extension of the experimental setup with respect to clustering (Supplementary Section B.5.4).
@@ B.5.2 Impact on prediction error estimates
For the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$, which is the focus of our illustration, this has two key implications: If $\widehat{\mathrm{PE}}_{\text{train}}$ results from an analysis setting where the apparent error was used to evaluate the final prediction model, the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ may underestimate the optimistic bias that would arise if $\mathcal{D_{\text{new}}}$ contained exclusively observations from new episodes and teams not present in $\mathcal{D}_{\text{train}}$. If $\widehat{\mathrm{PE}}_{\text{train}}$ corresponds to the simple or nested CV error, the clustering-induced optimistic bias would, under the assumption that $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ are subject to the same level of bias, effectively cancel out when considering the difference between $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$. However, as shown in Figure 5, this is not the case. Further analysis (not shown) reveals that the observed differences arise from the slightly higher proportion of patient episodes present in both $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ during resampling, compared to the proportion of episodes present in both $\mathcal{D}_{\text{train}}$ and $\mathcal{D_{\text{new}}}$ during the initial split. As a result, $\widehat{\mathrm{PE}}_{\text{train}}$ is affected by a larger optimistic bias than $\widehat{\mathrm{PE}}_{\text{new}}$, which manifests in Figure 5, where their difference is examined.
@@ B.5.4 Learning algorithms for clustered data
In addition to performing splits at the cluster level, we also extended the main experimental setup by including additional learning algorithms specifically designed for clustered data. These are the Random Effects/Expectation-Maximization Tree algorithm (REEMT; R package REEMtree; [128]), and the Linear Mixed-Effects Model Tree algorithm (LMMT; R package glmertree; [96]). In the implementation used for our illustration, both algorithms take into account the clustering structure by iterating between two steps: (i) fitting a decision tree using the CART algorithm for REEMT or the CIT algorithm for LMMT and (ii) estimating random intercepts via a linear mixed model, which are subtracted from the outcome variable in the subsequent tree-fitting iteration. To ensure model stability, random effects are only included for each palliative care team, rather than for each individual episode, as more than 300 episodes consist of only a single palliative care phase (Figure S7a). Including REEMT and LMMT in the analysis, however, does not yield new insights. Their results closely resemble those of CART and CIT, as demonstrated in Figure S9, which compares the prediction error differences of the algorithms.